Grade 12th passPhysical Chemistry

What is the importance of formula 2n^n?...........???

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Pawan Prajapati
5 Years ago
Explanation:

The 2n^2 rule is the first to be followed while filling up of electrons in the shells of an atom .

Name of shell | Shell number | 2n^2 |

K Shell | 1 | 2*1^1= 2

L Shell | 2 | 2*2^2=8

M Shell | 3 | 3*3^2=18

N Shell | 4 | 2*4^2=32

As seen above K Shell can a accommodate maximum 2 electrons as it is is the first Shell and closest to the nucleus of the atom . L Shell =8 and so on. The are two more rules to be followed while filling the shells of an electron .

2.Shells with lower energy will be filled with the maximum number of electrons first . Thereafter shells having higher energy get filled

3. The maximum number of electrons that can be contained in the otherwise Shell of the atom other wise known as the Valance Shell is 8 electrons .


Thanks